OpenStudy (anonymous):

-4.1k - 0.1k= 2.16-4.6k please heelp and show work so i can understand it further

pooja195 (pooja195):

−4.1k−0.1k=2.16−4.6k −4.1k+−0.1k=2.16+−4.6k (−4.1k+−0.1k)=−4.6k+2.16(Combine Like Terms) −4.2k=−4.6k+2.16 −4.2k=−4.6k+2.16 Any ideas what to do from here?

OpenStudy (anonymous):

its been a long time since ive done this from what i can see if i can remember correctly is to turn it positive right?

pooja195 (pooja195):

hmm not quite... we should add 4.6 to both sides to that we can easily isolate k after words so what is -4.2+4.6?

OpenStudy (anonymous):

2.4

pooja195 (pooja195):

Add 4.6k to both sides. −4.2k+4.6k=−4.6k+2.16+4.6k 0.4k=2.16 <----Not 2.4....its 0.4

pooja195 (pooja195):

Now divide 0.4 on both sides 0.4/0.4=1 so just k whats 2.16/0.4 ?

OpenStudy (anonymous):

oh my bad haha i didnt look at it right (x its uh 5.4? correct?

pooja195 (pooja195):

yep! :)

pooja195 (pooja195):

\(~~~~~~~~~~~~~~~~~~~~~~~~~~\Huge{\color{red}{\heartsuit}\color{orange}{\bigstar}\color{yellow}{\heartsuit}\color{green}{\bigstar}\color{blue}{\heartsuit}\color{purple}{\bigstar}\color{hotpink}{\heartsuit}}\)\(\Huge\bf\color{#FF0000}W\color{#FF4900}e\color{#FF9200}l\color{#FFDB00}c\color{#FFff00}o\color{#B6ff00}m\color{#6Dff00}e\color{#24ff00}~\color{#00ff00}t\color{#00ff49}o\color{#00ff92}~\color{#00ffDB}O\color{#00ffff}p\color{#00DBff}e\color{#0092ff}n\color{#0049ff}S\color{#0000ff}t\color{#2400ff}u\color{#6D00ff}d\color{#B600ff}y\color{#FF00ff}!\)\(~~~~~~~~~~~~~~~~~~~~~~~~~~\Huge{\color{red}{\heartsuit}\color{orange}{\bigstar}\color{yellow}{\heartsuit}\color{green}{\bigstar}\color{blue}{\heartsuit}\color{purple}{\bigstar}\color{hotpink}{\heartsuit}}\)\(\Large\bf\color{Lime}{Please,\ don't\ hesitate\ to\ ask\ questions!}\)\(\Large\bf\color{magenta}{We're\ here\ to\ help!}\Large\color{magenta}{\ddot\smile}\)

OpenStudy (anonymous):

awesome thank you thats the correct answer?

OpenStudy (anonymous):

and just double checking its 5.4k= 4.6+2.16+4.6